# Towards Resource Efficient and Interpretable Bias Mitigation in Large Language Models

### v1

@@ Abstract
Although large language models (LLMs) have demonstrated their effectiveness in a wide range of applications, they have also been observed to perpetuate unwanted biases present in the training data, potentially leading to harm for marginalized communities. In this paper, we mitigate bias by leveraging small biased and anti-biased expert models to obtain a debiasing signal that will be added to the LLM output at decoding-time. This approach combines resource efficiency with interpretability and can be optimized for mitigating specific types of bias, depending on the target use case. Experiments on mitigating gender, race, and religion biases show a reduction in bias on several local and global bias metrics while preserving language model performance.
1 Introduction
@@
Natural language generation (NLG) has risen in popularity greatly, serving as building blocks for applications such as chatbots, translators, and writing assistants and interacting with users in different domains [14]. One such example is OpenAI’s update in March 2021, which stated that GPT-3 powered applications generate an average of 4.5 billion words per day. However, despite recent advancements, these large language models (LLMs) have been reported to capture and reproduce unwanted biases and stereotypes [1, 21]. This occurs mainly because the large text corpora required for training such models are extracted from the Internet, which is not an accurate reflection of the diversity of real-world distributions. Generating biased outputs can result in serious negative consequences to society, ranging from offensive language that prevent certain demographic groups from adopting the technology [30] to biased job advertisements that discourage candidates from applying to certain positions [6].
@@
To address these issues, researchers have tried to curate better training data and improve the training process [10, 19]. Nevertheless, this approach remains less feasible in practice due to the significant human and computation resource involved. Recent works have instead focused on reducing the bias of generated outputs at decoding time to improve resource efficiency. For example, [30] introduced a prompt engineering method, Trigger, that concatenates a sequence of tokens to user inputs to reduce the output’s bias. However, the modified prompts suffer from a lack of interpretability and has been shown to spew racist output on non-racial contexts [35].
In this paper, we adapt the framework from [18]’s method of using small models as experts and anti-experts for detoxification. The biased and anti-biased experts in our system are small language models (LMs), for instance pre-trained GPT-2 Small models, fine-tuned on subsets of the RedditBias dataset [4]. They produce a debiasing signal that we incorporate into the LLM output at decoding-time. This approach combines resource efficiency - the experts require only several hundred sentences for fine-tuning, and interpretability - one can examine the shift in output probabilities for any given prompt. Furthermore, depending on the anticipated use case, our framework can be modified to address any type of bias expected.
@@
We evaluated the performance of the framework on three different bias directions, gender, race, and religion, and observed a reduction in bias on several local and global bias metrics. Since the experts rely on small biased datasets for fine-tuning, we substituted the fine-tuning dataset with StereoSet [23] and found that the results remain robust to the dataset choice; Stereotype Score was the only exception due to the data dependency of the bias metric. We also experimented with applying experts in one bias direction (race) to reduce bias for other directions (gender or religion). Doing so ascertains that optimizing for specific bias directions and use cases does not exacerbate the problem or create unwanted side effects. Last but not least, we investigated whether the debiased predictions follow human expectations by leveraging the framework’s interpretability and examining the debiasing signal and probability shifts. Comparing with Trigger, our results show similar levels of reduction in bias while better preserving overall LM performance, shedding deeper insight on the performance-fairness tradeoff. Our approach highlights the potential of decoding-time bias mitigation in handling real-world scenarios by providing resource efficiency and interpretability.
2 Related Work
@@
Recent research has shown that bias exists in many important natural language processing models, including word embeddings [5, 7, 13], question answering [25], and sentiment analysis [16]. NLG also suffers from similar problems as large language models are trained using text from the Internet [27], which likely contain unwanted stereotypes and skewed representations of true distributions. Bias in NLG negatively affect society in many ways, ranging from propagating and amplifying bias [1, 21] to discouraging certain groups from adopting the technology [30] or even increasing their vulnerability to harm and discrimination [32, 22]. However, despite recent efforts in measuring and creating benchmarks for bias in LLMs [15, 12], robustly quantifying the bias present remain difficult as traditional definitions of fairness [33] do not directly apply to the unstructured, open-ended text generated. In this paper, we consider a language generation model as biased if it disproportionately generates text that is often perceived as being negative, unfair, prejudiced, or stereotypical against an idea or a group of people with common attributes [9].
@@
There has been two main approaches to mitigating bias in NLG, namely domain-specific training and constrained decoding [37]. Domain-specific training requires fine-tuning the model with counterfactual data augmentation [10, 19], adversarial learning [20], or even regularization training [26]. Unfortunately, the excessive resource consumption of this approach makes it less feasible in real-world applications. The more resource-efficient approach focuses on guiding the output generation during decoding. For example, [30] mitigates bias by finding a sequence of tokens that, when concatenated to input prompts, will stimulate unbiased generation from the model - effectively modifying the user prompt before inference. However, the modified prompt may not be human readable and has been shown to spew racist output on non-racial contexts [35]. [17] proposed an Autoregressive Iterative Nullspace Projection algorithm that mitigates bias through a projection matrix that acts on the output logits. [28]’s method leverages the ability of a language model to recognize its own undesirable bias in a form of self-diagnosis. Last but not least, [37] considers detoxifying and debiasing a joint task, combining two adaptive approaches at inference time. Our approach incorporates both resource efficiency and interpretability: the experts are small language models and we can examine the shift in output probabilities before and after debiasing for any given prompt. Furthermore, our algorithm can be tailored to optimize for specific types of bias, depending on the anticipated use case of the generated text.
3 Methodology
3.1 Datasets
@@
Table 1: Sample counts and examples of data used to fine-tune experts and anti-experts for different bias directions
@@
Gender Stereotype 505 the cleaning lady is a smoking hot babe.
Anti-stereotype 505 the programming lady is probably cute and nice.
Race Stereotype 507 statistically black men are more dangerous.
@@
RedditBias is a real-world dataset for measuring and mitigating biases in dialog models. The corpus also demonstrates high levels of toxicity, which is a well-documented characteristic of Reddit discussions [3]. Even though bias in language systems can take many forms, we chose to adopt RedditBias in our analysis since it is publicly available and reflect common stereotypes directed towards historically non-dominant population groups [29]. The dataset contains pairs of targets T (demographic groups) and attributes A in the form of (T1, T2, A1, A2), where T1 denotes minority groups, T2 denotes dominant groups, and A1 and A2 the respective associated attributes with those groups. This allows us to generate fine-tuning data for stereotypes (T1, A1 and T2, A2) and anti-stereotypes (T1, A2 and T2, A1). Table 1 provides examples of the data used to fine-tune experts in our framework; note that only several hundred sentences are required, compared to the tens of millions needed to augment the LLM training process.
@@
We leverage BOLD prompts for evaluation. These prompts were extracted from English Wikipedia and aims to accurately reflect the diversity and structure of data given to text generation models. Their composition may influence the confidence of the evaluation but theoretically not the evaluation metric itself.
@@
We also use the StereoSet dataset for both fine-tuning and evaluation in this paper. The Intrasentence portion of the dataset consists of fill-in-the-blank sentences with three options. For example, the context sentence can be: Girls tend to be more BLANK than boys, followed by the stereotypical Option 1 soft, anti-stereotypical Option 2 determined, and unrelated Option 3 fish. The authors proposed two evaluation metrics, one for bias and one for performance, based on how the target LM chooses between the 3 options. Alternatively, when used for fine-tuning, one can generate stereotypical and anti-stereotypical data by completing the context sentence with the appropriate Option.
3.2 Bias Mitigation Framework
@@
Our framework leverages biased and anti-biased experts to incorporate a debiasing signal at decoding-time. To increase resource efficiency, the experts are small LMs, for example, GPT-2 Small, fine-tuned on small biased datasets. The expert, fine-tuned using anti-biased and anti-stereotypical data, represents a model with desirable attributes that overcome current societal stereotypes. Vice versa, the anti-expert is biased and reinforces current stereotypes. Intuitively, tokens considered both likely by the expert and unlikely by the anti-expert will have their probability increased at this intermediate output, known as the debiasing signal. By combining this interpretable debiasing signal with the target model (biased, not fine-tuned on any dataset), the framework generates less biased or unbiased outputs. Figure 1 illustrates how the framework incorporates an $\alpha$-weighted debiasing signal when given the prompt: The woman worked as a …, increasing the probability for the word doctor and decreasing that of nurse and babysitter. Compared to prior work that seeks to mitigate bias in the target model directly by improving the training data, the proposed framework is more resource efficient in terms of amount of data and computation power required through the usage of smaller expert models fine-tuned on small datasets.
@@
Mathematically, let us consider the case of conditional text generation with context $x_{1:t}=\{x_{1},\dots,x_{t}\}$. Let $\mathbf{z}_{t}\in R^{|V|}$ be the pre-softmax output of the target model, such that $P_{\theta}(x_{t}|x_{<t})=\mathrm{softmax}(\mathbf{z}_{t})$ is the probability distribution at step $t$. The goal of the approach is to modify $\mathbf{z}_{t}$ into $\tilde{\mathbf{z}}_{t}$, a debiased distribution that will promote unbiased text generation. Given the same context $x_{1:t}$, a forward pass through the expert (anti-biased model) gives the positive prediction $\mathbf{z}_{t}^{+}$, whereas the anti-expert (biased model) outputs $\mathbf{z}_{t}^{-}$. The algorithm combines these predictions with that of the original target LM in a way that promotes the most-likely tokens of the expert and demotes those of the anti-expert to reduce bias. It is important that the target model and the experts share the same vocabulary ${|V|}$, so that mathematical operations can be performed between $\mathbf{z}_{t},\mathbf{z}_{t}^{+},\mathbf{z}_{t}^{-}$. The resulting probability distribution resembles:
@@
In addition to resource efficiency and interpretability, the proposed approach also has the benefit of optimizing for anticipated use cases. Although we evaluate with experts fine-tuned on RedditBias, users can choose any available biased dataset in practice. For example, one can curate sentences pertaining to gender and occupations when using NLG for job advertisements. Since the majority of use cases revolve around mitigating bias for LLMs, we recommend fine-tuning datasets that consist of different types of bias representing current societal stereotypes.
4 Evaluation
4.1 Training Details and Resource Efficiency
We fine-tuned the pre-trained version of GPT-2 Small (124M parameters), the training is done on 2 epochs, learning rate of $10^{-5}$, batch size of 4, and Adam optimizer with $\beta_{1},\beta_{2}=(0.9,0.999)$, $\varepsilon=10^{-8}$. We performed a 90-10% train-validation split for all datasets described in Table 1. Our approach uses negligible resources compared to re-training a LLM; GPT-3 (175B parameters) would take 288 years on a single V100 NVIDIA GPU [24]. Fine-tuning the experts ranged from 107 to 483 seconds on V100 NVIDIA GPUs provided by the Satori IBM Power9 cluster.
4.2 Evaluation Metrics
@@
Global bias measures the difference of some high level property of the generated sentences between demographic groups. To generate the set of sentences needed for measuring global bias, we leveraged the BOLD dataset and generated 5 sentences per prompt. The Regard [31] metric aims to measure social perception and judgment towards the group. Since Regard captures the perception of the subject in the sentence, it serves as a better metric for evaluating bias than sentiment analysis, which captures the perception of the entire sentence, lower indicates less bias. The Toxicity metric focuses on the discrepancy in toxicity between the groups which we calculated using a RoBERTa classifier from [34], lower indicates less bias.
@@
4.3 Performance on Bias Mitigation
In this subsection, we analyze how the proposed bias mitigation framework performs with respect to the evaluation metrics. We adopted the default parameters of Hugging Face’s transformers library [36] using the decoding strategy of nucleus sampling with top-p=0.9, max-new-tokens=15, and a temperature of 1.0. We conducted studies with GPT-2 Medium as target models with the same expert settings of no debiasing (None), fine-tuning with aggregated data from all 3 bias directions (Full), and fine-tuning with only the specific bias direction (*bias). Tables 2 display results for gender and race bias (religion omitted for space); all experiments use a single run and regard, toxicity, and Hellinger distance has been scaled 100 times for easier observation.
Table 2: Debiasing results for gender and race bias with no debiasing (None), data from all bias directions (Full), and data only from that bias direction (Gender or Race). Best and second best results are indicated in bold and underlined, respectively. Arrows mark direction of highest performance, close to 50 is best for Stereotype Score SS.
Direction Debiasing Global bias Local bias Language Modeling
Regard $\downarrow$ Toxicity $\downarrow$ Hel. Dist. $\downarrow$ SS LM Score $\uparrow$ PPL $\downarrow$
Gender None 1.97 0.23 13.53 65.58 93.58 19.10
Full 1.47 0.18 12.98 63.12 92.40 20.12
Gender 2.07 0.31 13.27 65.94 93.11 19.36

### v2

@@ Abstract
Although large language models (LLMs) have demonstrated their effectiveness in a wide range of applications, they have also been observed to perpetuate unwanted biases present in the training data, potentially leading to harm for marginalized communities. In this paper, we mitigate bias by leveraging small biased and anti-biased expert models to obtain a debiasing signal that is added to the LLM output at decoding-time. This approach combines computational efficiency - fine-tuning a small model versus re-training a large model and interpretability - one can examine the probability shift from debiasing. The framework can also be tailored to specific contexts by switching the choice of the fine-tuning dataset. Experiments on mitigating gender, race, and religion biases on different architectures show a reduction in bias on several local and global bias metrics while preserving language model performance.
I Introduction
@@
Natural language generation (NLG) has become increasingly popular, serving as building blocks for applications such as chatbots, translators, and writing assistants and interacting with users in different domains [14]. However, despite recent advances, these large language models (LLMs) have been reported to capture and reproduce unwanted biases and stereotypes [1, 22]. This occurs mainly because the large text corpora required for training such models are extracted from the Web, which is not an accurate reflection of the diversity of real-world distributions. The generation of biased output can have serious negative consequences for society, ranging from offensive language that prevents certain demographic groups from adopting technology [31] to biased job ads that discourage candidates from applying to certain positions [6].
@@
To address these issues, researchers have tried to curate better training data and improve the training process [10, 20]. Nevertheless, this approach remains less feasible in practice due to the significant human and computation resources involved. Recent works have instead focused on reducing the bias of generated outputs at decoding time to improve computational efficiency. For example, [31] introduced a prompt engineering method, Trigger, which concatenates a sequence of tokens to user inputs to reduce the output’s bias. However, the modified prompts suffer from a lack of interpretability and the system has been shown to spew racist output on non-racial contexts [36].
In this paper, we adapt the detoxification methodology that uses small fine-tuned models [19] to mitigating bias. We use small language models (LMs), for instance pre-trained GPT-2 Small or LLaMA 3.2 1B models, fine-tuned on subsets of the RedditBias dataset [4], as the biased and anti-biased experts in our system. They produce a debiasing signal that is incorporated into the target LLM output at decoding-time; this LLM can have any architecture, provided it shares the same dictionary as the experts. This approach is computationally efficient because fine-tuning offers significant cost savings compared to re-training the entire model, and fine-tuning smaller expert models is more efficient than directly fine-tuning the target model. It is also interpretable, as one can examine the shift in output probabilities for any given prompt. Furthermore, if the context and cause of bias are known beforehand, the framework can be easily adjusted by fine-tuning with a targeted dataset.
@@
We evaluated the performance of the framework on two different target models along three different bias directions - gender, race, and religion. We observed a reduction in bias on several local and global bias metrics such as Regard [32], Toxicity, Hellinger distance, and Stereotype Score [24]. Since experts rely on small biased datasets for fine-tuning, we substituted the fine-tuning dataset with StereoSet [24]. We found that the results remain robust to the dataset choice; with the exception of the Stereotype Score due to the data dependency of the bias metric. We also experimented with applying experts in one bias direction (race) to reduce bias for other directions (gender or religion). Doing so ascertains that tailoring to specific bias directions and use cases does not exacerbate the problem or create unwanted side effects. Last but not least, we investigated whether the debiased predictions follow human expectations by leveraging the framework’s interpretability and examining the debiasing signal and probability shifts. Compared to the popular Trigger method, our results show similar levels of reduction in bias while preserving better overall LM performance and providing deeper insight on the performance-fairness tradeoff. Our approach highlights the potential of decoding-time bias mitigation in handling real-world scenarios by providing computational efficiency and interpretability.
II Related Work
@@
Recent research has shown that bias exists in many important natural language processing models, including word embeddings [5, 7, 13], question answering [26], and sentiment analysis [17]. NLG also suffers from similar problems as large language models are trained using text from the Internet [28], which likely contain unwanted stereotypes and skewed representations of true distributions. Bias in NLG negatively affect society in many ways, ranging from propagating and amplifying bias [1, 22] to discouraging certain groups from adopting the technology [31] or increasing their vulnerability to harm and discrimination [33, 23]. However, despite recent efforts in measuring and creating benchmarks for bias in LLMs [16, 12], robustly quantifying the bias present remain difficult as traditional definitions of fairness [34] do not directly apply to the unstructured, open-ended text generated. In this paper, we consider a language generation model as biased if it disproportionately generates text that is often perceived as being negative, unfair, prejudiced, or stereotypical against an idea or a group of people with common attributes [9].
@@
There has been two main approaches to mitigating bias in NLG, namely domain-specific training and constrained decoding [38]. Domain-specific training requires fine-tuning the model with counterfactual data augmentation [10, 20], adversarial learning [21], or even regularization training [27]. Unfortunately, the excessive resource consumption of this approach makes it less feasible in real-world applications. The more computationally efficient approach focuses on guiding the output generation during decoding. For example, [31] mitigates bias by finding a sequence of tokens that, when concatenated to input prompts, will stimulate unbiased generation from the model - effectively modifying the user prompt before inference. However, the modified prompt may not be human readable and has been shown to spew racist output on non-racial contexts [36]. [18] proposed an Autoregressive Iterative Nullspace Projection algorithm that mitigates bias through a projection matrix that acts on the output logits. [29]’s method leverages the ability of a language model to recognize its own undesirable bias in a form of self-diagnosis. Last but not least, [38] considers detoxifying and debiasing a joint task, combining two adaptive approaches at inference time. Our approach incorporates both computational efficiency and interpretability: the experts are small language models and we can examine the shift in output probabilities before and after debiasing for any given prompt.
III Methodology
We begin by describing the three datasets used throughout the experiments: the RedditBias dataset, the BOLD prompts [9], and the StereoSet dataset [24], followed by detailed descriptions of the bias mitigation framework.
III-A Datasets
@@
TABLE I: Sample counts and examples of data used to fine-tune experts and anti-experts for different bias directions
@@
Gender Stereotype 505 have it at my work the cleaning lady is a smoking hot babe.
Anti-stereotype 505 the programming lady is probably cute and really nice.
Race Stereotype 507 statistically straight black men are more dangerous.
@@
RedditBias is a real-world dataset for measuring and mitigating biases in dialog models. Although it proposes a dialog-oriented evaluation framework, which is considered too specific for the scope of this analysis, the data contains curated and high-quality (plausible and human-written) biased sentences extracted from Reddit threads. The corpus also demonstrates high levels of toxicity, which is a well-documented characteristic of Reddit discussions [3]. Even though bias in language systems can take many forms, we chose to adopt RedditBias in our analysis since it is publicly available and reflect common stereotypes directed towards historically non-dominant population groups [30]. The dataset contains pairs of targets T (demographic groups) and attributes A in the form of (T1, T2, A1, A2), where T1 denotes minority groups, T2 denotes dominant groups, and A1 and A2 the respective associated attributes with those groups. This allows us to generate fine-tuning data for stereotypes (T1, A1 and T2, A2) and anti-stereotypes (T1, A2 and T2, A1). Table I provides examples of the data used to fine-tune the experts in our framework; note that only several hundred sentences are required, compared to the tens of millions needed to augment the LLM training process.
@@
We include summary statistics for the BOLD prompts in Table II. These prompts were extracted from English Wikipedia and aims to accurately reflect the diversity and structure of data given to text generation models. Their composition may influence the confidence of the evaluation but theoretically not the evaluation metric itself.
@@
Table III shows the composition of the StereoSet dataset, used both for fine-tuning and evaluation in this paper. The Intrasentence portion of the dataset consists of fill-in-the-blank sentences with three options. For example, the context sentence can be: Girls tend to be more BLANK than boys, followed by the stereotypical Option 1 soft, anti-stereotypical Option 2 determined, and unrelated Option 3 fish. The authors proposed two evaluation metrics, one for bias and one for performance, based on how the target LM chooses between the 3 options. Alternatively, when used for fine-tuning, one can generate stereotypical and anti-stereotypical data by completing the context sentence with the appropriate Option.
TABLE II: BOLD statistics, we only considered gender, race, and religion in the evaluation.
Bias direction # of groups # of prompts
Profession 18 10,195
Gender 2 3,204
Race 4 7,657
Religion 7 639
Political ideology 12 1,984
TABLE III: Statistics of StereoSet dataset, Intrasentence CATS. Note that only Intrasentence CATs are considered when computing $S$ and LM Scores. This is because the Intersentence CAT requires adding a classification head on top of the model outputs, therefore modifying the architecture and potentially the evaluation.
Bias direction # Target # CATs Avg Len
Terms (triplets) (# words)
Gender 40 1,026 7.98
Profession 120 3,208 8.30
Race 149 3,996 7.63
Religion 12 623 8.18
Total 321 8,498 8.02
III-B Bias Mitigation Framework
@@
Our framework leverages biased and anti-biased experts to incorporate a debiasing signal at decoding-time. To increase computational efficiency, the experts are small LMs, for example, GPT-2 Small or LLaMA 3.2 1B, fine-tuned on small biased datasets. The expert, fine-tuned using anti-biased and anti-stereotypical data, represents a model with desirable attributes that overcome current societal stereotypes. Vice versa, the anti-expert is biased and reinforces current stereotypes. Intuitively, tokens considered both likely by the expert and unlikely by the anti-expert will have their probability increased at this intermediate output, known as the debiasing signal. By combining this interpretable debiasing signal with the target model (biased, not fine-tuned on any dataset), the framework generates less biased or unbiased outputs. Figure 1 illustrates how the framework incorporates an $\alpha$-weighted debiasing signal when given the prompt: The woman worked as a …, increasing the probability for the word doctor and decreasing that of nurse and babysitter. Compared to prior work that seeks to mitigate bias in the target model directly by improving the training data, the proposed framework is more computationally efficient in terms of amount of data and computation power required through the usage of smaller expert models fine-tuned on small datasets.
@@
Mathematically, let us consider the case of conditional text generation with context $x_{1:t}=\{x_{1},\dots,x_{t}\}$. Let $\mathbf{z}_{t}\in R^{|V|}$ be the pre-softmax output of the target model, such that $P_{\theta}(x_{t}|x_{<t})=\mathrm{softmax}(\mathbf{z}_{t})$ is the probability distribution at step $t$. The goal of the approach is to modify $\mathbf{z}_{t}$ into $\tilde{\mathbf{z}}_{t}$, a debiased distribution that will promote unbiased text generation. Given the same context $x_{1:t}$, a forward pass through the expert (anti-biased model) gives the positive prediction $\mathbf{z}_{t}^{+}$, whereas the anti-expert (biased model) outputs $\mathbf{z}_{t}^{-}$. The algorithm combines these predictions with that of the original target LM in a way that promotes the most-likely tokens of the expert and demotes those of the anti-expert to reduce bias. It is important that the target model and the experts share the same vocabulary |V|{|V|}, so that mathematical operations can be performed between $\mathbf{z}_{t},\mathbf{z}_{t}^{+},\mathbf{z}_{t}^{-}$. The resulting probability distribution resembles:
@@
In addition to computational efficiency and interpretability, the proposed approach also has the benefit of tailoring to specific contexts or bias directions. Although we evaluate with experts fine-tuned on RedditBias, users can choose any available examples and datasets of biased sentences in practice. For example, one can curate sentences pertaining to gender and occupations when using NLG for job advertisements.
Compared to directly fine-tuning the target model with anti-stereotypical data, which is also much more efficient than re-training the LLM, our approach still offers some computation savings by fine-tuning a smaller version of the same architecture. Furthermore, the scaling parameter $\alpha$ can control the strength of the debiasing signal and one can examine the signal to observe the probability shift, while retaining the original model outputs as reference for interpretability.
IV Evaluation
We present training details for the experts and evaluation metrics before showing mitigation results for the different directions of bias. We also performed in-depth analysis on the effects of different fine-tuning datasets, interactions between bias directions, and interpretations of the debiasing signal.
IV-A Training Details and Computational Efficiency
We fine-tuned the pre-trained version of GPT-2 Small (124M parameters) and LLaMA 3.2 1B (1.23B parameters) to create the expert models. The training is done on 2 epochs, learning rate of $10^{-5}$, batch size of 4, and Adam optimizer with $\beta_{1},\beta_{2}=(0.9,0.999)$, $\varepsilon=10^{-8}$. We performed a 90-10% train-validation split for all datasets described in Table I.
Our approach uses negligible time compared to re-training an LLM; GPT-3 (175B parameters) would take 288 years on a single V100 NVIDIA GPU [25]. Fine-tuning the experts took around 5 minutes on V100 NVIDIA GPUs provided by the Satori IBM Power9 cluster. For comparison, directly fine-tuning the target models, in our experiments GPT-2 Medium (345M parameters) and LLaMA 3.2 3B (3.21B parameters), took around 20 minutes; we ran out of memory when fine-tuning LLaMA and applied Low-Rank Adaptation (LoRA) [15]. In general, the approach is much more computationally efficient than retraining the LLM and more efficient than directly fine-tuning the target model, especially for the larger versions like LLaMA 3.2 90B.
IV-B Evaluation Metrics
@@
Global bias measures the difference of some high level property of the generated sentences between demographic groups. To generate the set of sentences needed for measuring global bias, we leveraged the BOLD dataset and generated 5 sentences per prompt. The Regard [32] metric aims to measure social perception and judgment towards the group. Since Regard captures the perception of the subject in the sentence, it serves as a better metric for evaluating bias than sentiment analysis, which captures the perception of the entire sentence. The implementation consists of a probabilistic text classifier trained on 1700 human-labeled samples of biased language and computes the social perception and judgement (positive, neutral, and negative) towards the group. We then compute the difference in Regard between groups, lower indicates less bias. The Toxicity metric focuses on the discrepancy in toxicity between the groups which we calculated using a RoBERTa classifier from [35], lower indicates less bias.
@@
IV-C Adjusting the Strength of the Debiasing Signal
The $\alpha$ hyperparameter determines the importance of the debiasing signal relative to the target model output. In order to determine the optimal hyperparameter value, we experimented with values between 0 to 5 and evaluated against the Hellinger Distance, SS, and PPL metrics. As expected, LM performance decreases as $\alpha$ increases. The Hellinger Distance is lowest when $\alpha$ is 3 and the SS score is closest to 50% when $\alpha$ is 1. Thus, all the experiments were conducted with a weight of 1.0 for the best performance-fairness tradeoff. Directly fine-tuning the target model loses this important scaling factor: Adjusting the number of epochs only provides an inaccurate estimation.
IV-D Performance on Bias Mitigation
In this subsection, we analyze how the proposed bias mitigation framework performs with respect to the evaluation metrics. We adopted the default parameters of Hugging Face’s transformers library [37] using the decoding strategy of nucleus sampling with top-p=0.9, max-new-tokens=15, and a temperature of 1.0. We conducted experiments using GPT-2 Medium as target model with GPT-2 Small experts and LLaMA 3.2 3B as target model with LLaMA 3.2 1B experts. For both models, we compared performance-fairness tradeoffs for models with no debiasing (None), full framework with expert and anti-expert (Proposed), the anti-expert only setting (Anti-only), and directly fine-tuning the target model (Fine-tuned). Tables IV, V, and VI display results for gender, race, and religion bias respectively; all experiments use a single run and regard, toxicity, and Hellinger distance has been scaled 100 times for easier observation.
TABLE IV: Debiasing results for gender bias with no debiasing (None), full framework with expert and anti-expert (Proposed), anti-expert only setting (Anti-only), and directly fine-tuning the target model (Fine-tuned). Best and second best results are indicated in bold and underlined, respectively. Arrows mark direction of highest performance, close to 50 is best for Stereotype Score $S$.
Target Model Debiasing Global bias Local bias Language Modeling
Regard $\downarrow$ Toxicity $\downarrow$ Hel. Dist. $\downarrow$ $S$ LM Score $\uparrow$ PPL $\downarrow$
GPT-2 Medium None 1.97 0.23 13.53 65.58 93.58 19.10
Finetuned 3.06 2.66 13.44 62.45 93.05 27.12
Proposed 1.47 0.18 12.98 63.12 92.40 20.12
Anti-only 0.85 0.09 15.48 65.44 90.60 27.06